Solve du/dt =DT/


1
Expert's answer
2021-04-29T14:47:29-0400

"du = cv.dt" (for an ideal gas)


"\dfrac{du}{dt} = cv"


"\dfrac{du}{dt} = DT/DU × D²U/D²T"


"\dfrac{dU}{dT }= cv"